# Grasping the Essentials: Tailoring Large Language Models for Zero-Shot Relation Extraction

### v1

@@ Appendix G Case Study
In order to showcase how LLMs function in accordance with our design expectations, we provide some dialogue cases in Fig. 8 and Fig. 9. From the example dialogues, we can see that GPT-4 is considerably capable of identifying and summarizing the existing patterns in the dialogue history. The conversational feature also improves the interpretability of the results which can be further exploited with human-in-the-loop methods for future research or downstream applications.
We can also see that feedback-driven negative relation definition generation is of great potential. As the feedback instances may contain false predictions which can be directly taken by LLM to generate new negative relations and effectively recitify the SLM’s bias. Such feature directly boost the precision score as shown in Sec. 5.2. Hence, we can see that our design expectations are well fulfilled and this also qualitatively explains the performance boost of our proposed model.
Appendix H Error Analysis
  P40: <ENT1> was/is the child (not stepchild) of <ENT0>
1. <ENT1> Lord Louis Mountbatten </ENT1> ( Hugh Bonneville ) arrives at Viceroy ’s House in Delhi in 1947 with his strong - willed wife <ENT0> Edwina </ENT0> ( Gillian Anderson ) and daughter Pamela . (Gold: P26 | Pred: P40 | Pos Prob: 0.995)
2. Daughter of <ENT1> Sancho IV </ENT1> and of <ENT0> María de Molina </ENT0> , Infanta Beatrice was born in Toro . (Gold: P26 | Pred: P40 | Pos Prob: 0.998)
3. <ENT1> Ruby Aldridge </ENT1> is the daughter of former Playboy playmate Laura Lyons and artist and graphic designer Alan Aldridge , and younger sister of fashion model <ENT0> Lily Aldridge </ENT0> . (Gold: P3373 | Pred: P40 | Pos Prob: 0.959)
4. <ENT0> Jan van Hoogstraten </ENT0> ( 1628–1654 ) was a Dutch Golden Age painter who died young in Vienna , where he was living with his older brother <ENT1> Samuel van Hoogstraten </ENT1> . (Gold: P3373 | Pred: P40 | Pos Prob: 0.998)
1. In November 1966 , retired <ENT1> USMC </ENT1> Major <ENT0> Donald Keyhoe </ENT0> and Richard H. Hall , both of NICAP , briefed the panel . (Gold: P241 | Pred: P410 | Pos Prob: 0.996)
2. General Sir <ENT0> Hew Whitefoord Dalrymple </ENT0> , 1st Baronet ( 3 December 1750 – 9 April 1830 ) was a <ENT1> British Army </ENT1> general and Governor of Gibraltar . (Gold: P241 | Pred: P410 | Pos Prob: 0.995)
P57: <ENT1> was/is the director(s) of <ENT0> (a film, TV-series, stageplay, video game or similar)
P84 (131): <ENT1> was/is the architect or architectural firm responsible for designing <ENT0> (a building)
1. The <ENT0> Yokosuka Arts Theatre </ENT0> , part of the Bay Square complex by <ENT1> Kenzo Tange </ENT1> , is a venue for opera , orchestral concerts , chamber music , and films . (Gold: P84 | Pred: P57 | Pos Prob: 0.981)
2. Memorials such as <ENT1> Maya Lin </ENT1> ’s <ENT0> Vietnam Veterans Memorial </ENT0> and Peter Eisenman ’s Memorial to the Murdered Jews of Europe are also said to reflect themes of trace and erasure . (Gold: P84 | Pred: P57 | Pos Prob: 0.979)
Table 12 shows the major source of false positive predictions of the final tuned SLM RE model, it can be seen that the majority of false positive predictions are concentrated on a few similar negative relations. As our proposed task setting assumes the unknown negative relation space, one challenge appears to be learning the positive relation against the unknown and infinitely many negative relations. To address such challenge, our model derive targeted negative relations based on the feedback of model inference. Based on the results in Table 12, we can see that the challenge is not fully eliminated which serves as a promising future research direction. Another feature seen from the false positive predicted instances is that some typical false positive instances actually express the target relation in addition to its gold relation. However, the target relation is not expressed by the tagged entity mention pair. This may indicate that better RE architectures which well model the position awareness of target entity pairs can be adopted for improving the overall performance. We also leave this for future studies.

### v2

@@ Appendix I Case Study
In order to showcase how LLMs function in accordance with our design expectations, we provide some dialogue cases in Fig. 8 and Fig. 9. From the example dialogues, we can see that GPT-4o is considerably capable of identifying and summarizing the existing patterns in the dialogue history. The conversational feature also improves the interpretability of the results which can be further exploited with human-in-the-loop methods for future research or downstream applications.
It’s also evident that feedback-driven negative relation definition generation is of great potential. As the feedback instances may contain false predictions which can be directly taken by LLM to generate new negative relations and effectively rectify the SLM’s bias. Such feature directly boost the precision score as shown in Sec. 5.2. Hence, we can see that our design expectations are well fulfilled and this also qualitatively explains the performance boost of our proposed model.
Meanwhile, Fig. 9 also indicates that GPT-4o is still not perfect on giving the correct intermediate analysis of examining each of the feedback example. This indicates the future work can further improve the design of feedback acquisition and feedback processing.